# Towards One-step Causal Video Generation via Adversarial Self-Distillation

### v1

@@ Appendix B Hyperparameter Analysis
To investigate the impact of the ASD loss on model performance, we evaluated different variants of the hyperparameter $\alpha$ in the Equation˜5 and computed the corresponding Total Score on VBench under three inference settings, as illustrated in Figure˜8. The results indicate that at 4-step generation, the Total Scores across different model variants are comparable. When using $2^{*}$-step generation, the variant without ASD Loss ($\alpha=0$) exhibits a noticeable performance drop, while other variants maintain performance levels similar to those observed in 4-step generation. Under $1^{*}$-step generation, all methods experience a decline in performance, with the most pronounced degradation occurring when $\alpha$ is set to zero. These findings demonstrate that the proposed ASD training objective enhances video quality under skip generation settings and remains robust across a wide range of hyperparameter values.
Appendix C User Study Details
We randomly selected 20 prompts from the VBench(Huang et al., 2024b) prompts expanded by Self-Forcing. Five baseline methods Sand-AI (2025); Huang et al. (2025); Yin et al. (2025); Chen et al. (2025); Wang et al. (2025) were used to generate corresponding videos, which were then paired with corresponding videos generated by our method, resulting in $5\times 20$ video pairs. During the user study, each participant evaluated 20 video pairs, including 4 randomly selected pairs from each baseline. The final results were averaged across all participants. The user study interface, as shown in Figure˜9, displayed a video pair alongside the corresponding prompt and selection buttons. Participants could choose whether the left video was better, the right video was better, or both were of similar quality. To ensure a fair comparison with both one-step and two-step configurations, we retrained the one-step and two-step versions of Self-Forcing.
@@ Appendix C User Study Details
Appendix D Usage of LLMs

### v2

@@ Appendix F Hyperparameter Analysis
To investigate the impact of the ASD loss on model performance, we evaluated different variants of the hyperparameter $\alpha$ in the Equation˜5 and computed the corresponding Total Score on VBench under three inference settings, as illustrated in Figure˜11. The results indicate that at 4-step generation, the Total Scores across different model variants are comparable. When using $2^{*}$-step generation, the variant without ASD Loss ($\alpha=0$) exhibits a noticeable performance drop, while other variants maintain performance levels similar to those observed in 4-step generation. Under $1^{*}$-step generation, all methods experience a decline in performance, with the most pronounced degradation occurring when $\alpha$ is set to zero. These findings demonstrate that the proposed ASD training objective enhances video quality under skip generation settings and remains robust across a wide range of hyperparameter values. While absolute values vary slightly across $\alpha$, the trend is consistent: skip-step performance improves markedly whenever $\alpha\neq 0$.
Appendix G User Study Details
We randomly selected 20 prompts from the VBench(Huang et al., 2024b) prompts expanded by Self-Forcing. Five baseline methods Sand-AI (2025); Huang et al. (2025); Yin et al. (2025); Chen et al. (2025); Wang et al. (2025) were used to generate corresponding videos, which were then paired with corresponding videos generated by our method, resulting in $5\times 20$ video pairs. During the user study, each participant (a total of 12 participants) evaluated 20 video pairs, including 4 randomly selected pairs from each baseline. The order and pairing of videos were independently randomized for every participant. Participants could not see each other’s choices. The final results were averaged across all participants. The user study interface, as shown in Figure˜12, displayed a video pair alongside the corresponding prompt and selection buttons. Participants could choose whether the left video was better, the right video was better, or both were of similar quality. To ensure a fair comparison with both one-step and two-step configurations, we retrained the one-step and two-step versions of Self-Forcing.
@@ Appendix G User Study Details
Appendix H Usage of LLMs
@@
(a) Self Forcing
(b) Ours
Figure 13: The DMD loss curves of Self Forcing and our method during training. The DMD loss values remain around 0.2 for both methods, but our method exhibits substantially lower fluctuation.
Appendix I Training Dynamics Analysis
Figure 14: The discriminator’s predicted logit difference between n-step and (n+1)-step samples during training.
Figure˜13 presents the DMD loss curves of our method and Self-Forcing during training. Notably, our approach exhibits significantly reduced fluctuations in DMD loss compared to Self-Forcing. With ASD, the DMD loss shows lower mean (0.180 vs. 0.198) and dramatically reduced variance (0.00258 vs. 0.01963) compared to pure DMD training, demonstrating that the proposed ASD loss effectively stabilizes distribution matching with the teacher model during training.
Figure˜14 shows that the discriminator’s logit difference between n-step and (n+1)-step samples oscillates calmly around zero, indicating that the generator and discriminator co-evolve stably rather than collapsing or diverging. This behavior supports that ASD training is stable in practice, despite using an adversarial term.
Training Overhead. The ASD framework introduces only a lightweight discriminator on top of the existing TA score model’s backbone, using shared parameters for all n-step discriminators. This results in minimal extra memory usage compared to traditional DMD. The additional GPU memory required by ASD is +1.3% compared to the Self-Forcing baseline, which is a modest increase. The additional training time is approximately +20% relative to Self-Forcing. Given that ASD allows a single model to handle multiple inference steps (1, 2, 3, 4), this overhead is relatively small compared to the computational cost of training separate models for each step.